# Implicit Regularization of Infinitesimally-perturbed Gradient Descent Toward Low-dimensional Solutions

### v1

@@ Abstract
Implicit regularization refers to the phenomenon where local search algorithms converge to low-dimensional solutions, even when such structures are neither explicitly specified nor encoded in the optimization problem. While widely observed, this phenomenon remains theoretically underexplored, particularly in modern over-parameterized problems. In this paper, we study the conditions that enable implicit regularization by investigating when gradient-based methods converge to second-order stationary points (SOSPs) within an implicit low-dimensional region of a smooth, possibly nonconvex function. We show that successful implicit regularization hinges on two key conditions: $(i)$ the ability to efficiently escape strict saddle points, while $(ii)$ maintaining proximity to the implicit region. Existing analyses enabling the convergence of gradient descent (GD) to SOSPs often rely on injecting large perturbations to escape strict saddle points. However, this comes at the cost of deviating from the implicit region. The central premise of this paper is that it is possible to achieve the best of both worlds: efficiently escaping strict saddle points using infinitesimal perturbations, while controlling deviation from the implicit region via a small deviation rate. We show that infinitesimally perturbed gradient descent (IPGD), which can be interpreted as GD with inherent “round-off errors”, can provably satisfy both conditions. We apply our framework to the problem of over-parameterized matrix sensing, where we establish formal guarantees for the implicit regularization behavior of IPGD. We further demonstrate through extensive experiments that these insights extend to a broader class of learning problems.
keywords:
Nonconvex optimization, Gradient descent, Implicit regularization
@@ 1 Introduction
Over-parameterization, where the sheer number of variables surpasses what is essential for identifying the optimal solution, is a ubiquitous phenomenon in contemporary optimization problems in machine learning (ML). In the face of over-parameterization, gradient descent (GD) or its variants may encounter significant statistical gaps as they become prone to overfitting. However, recent research has postulated that these algorithms exhibit an implicit regularization in a wide range of optimization problems—spanning from the classical sparse recovery problem [1, 2, 3, 4] to low-rank matrix and tensor optimization [5, 6, 7, 8, 9] to modern neural network training [10, 11, 12, 13, 14, 15]—favoring solutions with low dimensionality, even when such low-dimensional structures are neither explicitly characterized nor encoded in the problem formulation. This property mitigates overfitting and enhances the interpretability of the solutions. Examples of implicit regularization of GD are pervasive across a wide range of learning problems:
  • Sparse recovery: In sparse recovery problem, the goal is to obtain a sparse parameter vector $\bm{\theta}^{\star}\in\mathbb{R}^{d}$ with at most $s\ll d$ nonzero entries that minimizes an empirical loss function $L(\bm{\theta})$. Recent observations have revealed that, when the unknown parameter is represented as a two-layer model $\bm{\theta}=\mathbf{u}\odot\mathbf{u}-\mathbf{v}\odot\mathbf{v}$ (where the notation $\odot$ denotes the element-wise product), GD with a small initial point applied to the lifted objective $f(\mathbf{u},\mathbf{v}):=L(\mathbf{u}\odot\mathbf{u}-\mathbf{v}\odot\mathbf{v})$ converges to the true $s$-sparse parameter $\bm{\theta}^{\star}$ [1, 4], even though denser solutions may lead to the same or smaller loss values. Remarkably, this result holds even when the sparsity parameter $s$ is entirely unknown or unspecified. [^1] A similar result has been established for a slightly different model, $\bm{\theta}=\mathbf{u}\odot\mathbf{v}$ [2, 3].
  • Low-rank matrix recovery: Another class of problems for which GD exhibits a natural implicit regularization is low-rank matrix recovery. A prominent example is the matrix sensing problem, where the goal is to recover an $n\times n$ positive semi-definite (PSD) matrix ${\mathbf{M}^{\star}}$ of rank $r$ from a limited number of Gaussian measurements by minimizing a loss function $L({\mathbf{M}^{\star}})$. A common approach involves factorizing ${\mathbf{M}^{\star}}$ as $\mathbf{X}=\mathbf{X}\mathbf{X}^{\top}$ and directly optimizing the lifted loss $f(\mathbf{X})=L(\mathbf{X}\mathbf{X}^{\top})$ over the $n\times r^{\prime}$ factor matrix $\mathbf{X}$, where $r^{\prime}\geq r$. Similar to the $s$-sparse recovery, GD with a small initial point applied to $L(\mathbf{X}\mathbf{X}^{\top})$ converges to the rank-$r$ matrix ${\mathbf{M}^{\star}}$, even when $r$ is unknown and severely over-estimated, up to $r^{\prime}=n$ [5, 6, 16].
  • Models with sparse basis function decomposition: A broader class of learning problems where the implicit regularization of GD toward low-dimensional solutions is empirically observed is those with sparse basis function decomposition, which encompasses both the sparse recovery and low-rank matrix recovery problems discussed above, as well as tensor decomposition and various types of nonlinear neural networks [17]. Further details on these models are provided later in the paper.
Despite these intriguing and promising findings, the theoretical foundations of the implicit regularization induced by GD in over-parameterized settings remain largely underexplored and are often limited to specific, idealized scenarios. In this paper, we aim to address this gap by presenting unified conditions that enable the implicit regularization in GD, potentially extending its applicability to broader classes of over-parameterized optimization problems.
More concretely, we consider a twice-differentiable, potentially nonconvex function $f:\mathbb{R}^{d}\to\mathbb{R}$ endowed with an implicit region $\mathcal{M}\subset\mathbb{R}^{d}$, presumed to possess a low dimension, specifically $\dim(\mathcal{M})=k\ll d$, capturing the reduced dimensionality of the desired solution.
While the existence of a low-dimensional implicit region can often be justified and anticipated in various learning tasks, its explicit characterization is task-dependent and typically unknown a priori. For example, for the aforementioned sparse recovery problem, this implicit region can be defined as =$\mathcal{M}=\{(\mathbf{u},\mathbf{v}):\mathbf{u}_{i}=\mathbf{v}_{i}=0\ \text{if}\ \bm{\theta}^{\star}_{i}=0\}$ with $\dim(\mathcal{M})=2s$, which depends on the unknown sparsity pattern of the $s$-sparse solution $\bm{\theta}^{\star}$. Similarly, for the low-rank matrix recovery problem, a possible choice of the implicit region is =$\mathcal{M}=\{\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}:\mathrm{rank}(\mathbf{X})\leq r\}$ with $\dim(\mathcal{M})=O(dr)$, which again depends on the unknown rank $r$. While conventional wisdom advocates for the use of explicit regularization or constraints to promote low-dimensional structure in the solution, recent findings suggest that, in both problems, GD with a small initialization tends to remain close to these (unknown) implicit regions.
Motivated by these intriguing observations, our goal is to identify general conditions on the loss function $f$ and the implicit region $\mathcal{M}$ under which GD converges to a second-order stationary point (SOSP) of $f$ within $\mathcal{M}$, referred to as an $\mathcal{M}$-SOSP and mathematically characterized as:
@@
Our interest in an $\mathcal{M}$-SOSP is motivated by the key observation that, for generic functions, any $\mathcal{M}$-SOSP corresponds to a local minimum of $f$ within $\mathcal{M}$ [18, Theorem 2.9], which, by definition of $\mathcal{M}$, is low-dimensional. From a learning perspective, the low-dimensional nature of the local solutions within the implicit region $\mathcal{M}$ promotes better generalization and enhances interpretability. In contrast, local solutions outside the implicit region $\mathcal{M}$ tend to be less interpretable and more susceptible to overfitting.
Evidently, the successful convergence of GD to $\mathcal{M}$-SOSPs critically depends on both the structure of the implicit region $\mathcal{M}$ and the properties of the loss function $f$. In this work, we identify two key factors governing this convergence: $(i)$ the ability to efficiently escape strict saddle points, which are points that are first-order but not second-order stationary, while $(ii)$ maintaining proximity to the implicit region $\mathcal{M}$. As will be explained next, the existing analyses that enables the convergence of GD to SOSPs often rely on applying large perturbations to the iterates to ensure an efficient escape from strict saddle points, but this typically comes at the cost of significantly deviating from $\mathcal{M}$. The main premise of this paper is that it is possible to achieve the best of both worlds: efficiently escaping from strict saddle points via infinitesimal perturbations, while controlling the deviation from $\mathcal{M}$ by quantifying the deviation rate of $f$ from $\mathcal{M}$. We prove that the implicit regularization of GD toward an $\mathcal{M}$-SOSP is achieved when the algorithm leverages both infinitesimal perturbations and a small deviation rate.
1.1 A Major Theoretical Hurdle in Over-parameterized Regime
It has been shown that various gradient-based algorithms—such as perturbed gradient descent (PGD) [19, 20] and stochastic gradient descent (SGD) [21]—can escape strict saddle points and converge to an approximate SOSP as effectively as their second-order counterparts.
@@ 1.1 A Major Theoretical Hurdle in Over-parameterized Regime
To illustrate the underlying success mechanisms of these gradient-based algorithms, let us consider PGD. Starting from an initial point $\mathbf{x}_{0}$, PGD iteratively updates the solution using the rule $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\nabla f(\mathbf{x}_{t})+\bm{\xi}_{t}$, where $\bm{\xi}_{t}$ represents an isotropic perturbation uniformly drawn from the Euclidean ball with radius $\gamma$ centered at zero, $\mathbb{B}(\gamma)$, when $\left\lVert\nabla f(\mathbf{x}_{t})\right\rVert$ is sufficiently small; otherwise, $\bm{\xi}_{t}$ is set to zero [20, Algorithm 1]. The reasoning behind this episodal perturbation is natural: when $\left\lVert\nabla f(\mathbf{x}_{t})\right\rVert$ is small, the iterate $\mathbf{x}_{t}$ is either near a strict saddle point or an SOSP. Unlike SOSPs, strict saddle points are unstable fixed points of GD [22, Theorem III.7]. Consequently, an isotropic random perturbation of sufficient magnitude enables GD to escape from these strict saddle points efficiently [20, Theorem 3]. We note that a similar argument applies to a non-periodic variant of PGD [23, Theorem 4.1], and SGD under the dispersive noise assumption for the gradients [21, Assumption 4], as well as the earlier approaches such as noisy stochastic GD [19, Algorithm 1], and saddle-normalized GD [24, Algorithm 1]. Moreover, such additive perturbations appear necessary, as without them, vanilla GD may take exponential time to escape strict saddle points [25, Theorem 4.1].
@@
Figure 1: PGD applied to matrix sensing converges at a linear rate under exact parameterization but fails to converge with over-parameterization. In this example, the true matrix ${\mathbf{M}^{\star}}$ is a $20\times 20$ PSD matrix with rank $r=3$. The search rank is set to $r^{\prime}=3$ and $r^{\prime}=4$ for exact- and over-parameterized settings.
@@
Figure 2: When perturbation is added to the gradient, it contributes to two directions: deviation direction, which steers the iterations away from the implicit region $\mathcal{M}$, and the negative Hessian direction, which enables an escape from strict saddle points.
@@
Unfortunately, in over-parameterized settings, guaranteed escape from strict saddle points alone is insufficient to ensure convergence to an $\mathcal{M}$-SOSP. In fact, even a slight over-parameterization can disrupt the convergence of PGD, potentially causing it to fail. Figure 2 shows the performance of PGD on an instance of matrix sensing (introduced in Section 1). When the rank $r$ of the target matrix is known, PGD with a random initial point applied to $f(\mathbf{X})=L(\mathbf{X}\mathbf{X}^{\top})$ with $\mathbf{X}\in\mathbb{R}^{n\times r}$ converges to the ground truth matrix at a linear rate, without getting trapped at a strict saddle point. However, even a slight over-parameterization, with a search rank of $r^{\prime}=r+1$, causes PGD to converge to a solution with a rank strictly greater than $r$, thereby incurring significant errors.
@@
The failure of PGD to converge to an $\mathcal{M}$-SOSP in the over-parameterized regime is precisely due to the impact of perturbations, as shown in Figure 2. When an isotropic random perturbation is applied, it inevitably affects two directions: a direction spanned by the eigenvectors corresponding to the negative eigenvalues of the Hessian, which pushes the iterate away from the strict saddle point (shown as $(a)$), and another direction that causes the iterate to deviate away from the implicit region $\mathcal{M}$ (shown as $(b)$). When the perturbation is both isotropic and large, these directions become comparable, causing the PGD iterates to escape the saddle point but simultaneously drift away from $\mathcal{M}$, ultimately leading to the failure of the algorithm.

### v2

# Understanding the Implicit Regularization of Gradient Descent in Over-parameterized Models

@@ Abstract
Implicit regularization refers to the tendency of local search algorithms to converge to low-dimensional solutions, even when such structures are not explicitly enforced. Despite its ubiquity, the mechanism underlying this behavior remains poorly understood, particularly in over-parameterized settings. We analyze gradient descent dynamics and identify three conditions under which it converges to second-order stationary points within an implicit low-dimensional region: (i) suitable initialization, (ii) efficient escape from saddle points, and (iii) sustained proximity to the region. We show that these can be achieved through infinitesimal perturbations and a small deviation rate. Building on this, we introduce Infinitesimally Perturbed Gradient Descent (IPGD), which satisfies these conditions under mild assumptions. We provide theoretical guarantees for IPGD in over-parameterized matrix sensing and empirical evidence of its broader applicability.
@@ 1 Introduction
Over-parameterization, where the sheer number of variables surpasses what is essential for identifying the optimal solution, is a ubiquitous phenomenon in contemporary optimization problems in machine learning (ML). In the face of over-parameterization, gradient descent (GD) or its variants may encounter significant statistical gaps as they become prone to overfitting. However, recent research has postulated that these algorithms exhibit an implicit regularization in a wide range of problems—spanning from sparse recovery problem [vaskevicius2019implicit, zhao2019implicit, ma2022blessing, woodworth2020kernel] to low-rank matrix and tensor optimization [li2018algorithmic, stoger2021small, soltanolkotabi2023implicit, wang2020beyond, tong2022scaling] to neural network training [belkin2019reconciling, yang2020rethinking, arora2018optimization, huh2021low, neyshabur2014search, neyshabur2017implicit]—favoring solutions with low dimensionality, even when such low-dimensional structures are neither explicitly specified nor encoded in the problem.
More specifically, consider the general task of estimating unknown parameters $\bm{\theta}$ of a model by minimizing a loss function $L:\mathbb{R}^{n}\to\mathbb{R}$:
$\min_{\bm{\theta}\in\mathbb{R}^{n}}\ L(\bm{\theta})\qquad\text{subject to}\qquad\bm{\theta}\in\mathcal{D}.$ (1)
In the above optimization, $\mathcal{D}\subseteq\mathbb{R}^{n}$ denotes the feasible region containing the underlying ground truth parameter $\bm{\theta}^{\star}$ of the model, frequently presumed to possess a low dimension, specifically $\dim(\mathcal{D})=k\ll n$. Many learning tasks are naturally endowed with such low-dimensional structures: Consider the following examples:
  • Sparse recovery: In the sparse recovery problem, the goal is to obtain a sparse vector $\bm{\theta}^{\star}\in\mathbb{R}^{n}$ with at most $k\ll n$ nonzero entries that minimizes a loss function $L(\bm{\theta})$ [mazumder2023subset, hazimeh2020fast]. In this case, the feasible region can be defined as $\mathcal{D}=\{\bm{\theta}\in\mathbb{R}^{n}:\bm{\theta}_{i}=0,\ \text{if $\bm{\theta}^{\star}_{i}=0$}\}$, which satisfies $\dim(\mathcal{D})=k$.
  • Low-rank matrix/tensor recovery: Another category of problems that reveal naturally low-dimensional structures is found in low-rank matrix recovery [chi2019nonconvex]. A prominent example is the matrix sensing problem, where the goal is to recover an $n\times n$ positive semi-definite (PSD) matrix of rank at most $r$ from a limited number of random measurements. In this setting, the feasible region is $\mathcal{D}=\{\mathbf{\Theta}\in\mathbb{R}^{n\times n}:\mathbf{\Theta}\succeq 0,\mathrm{rank}(\mathbf{\Theta})\leq r\}$ with $\dim(\mathcal{D})=\mathcal{O}(nr)$, where $\mathbf{\Theta}\succeq 0$ imposes PSD constraint. This setting can be naturally extended to low-rank tensor recovery, where the goal is to recover an order-$L$ tensor of rank $r$ [cai2019nonconvex]. In this scenario, the feasible region $\mathcal{D}$ can be defined as the set of order-$L$ tensors of rank $r$.
A critical challenge in solving (1), however, is that the feasible region $\mathcal{D}$ is rarely known in practice. For example, in sparse recovery, it is unreasonable to assume prior knowledge of the support of the true parameter $\bm{\theta}^{\star}$, which is crucial in characterizing $\mathcal{D}$. Similarly, in low-rank models, the true rank $r$ is rarely known a priori. A common workaround is to introduce a regularizer that promotes the low-dimensional structure underlying $\mathcal{D}$. However, this strategy requires tuning regularization parameters (e.g., chosen via cross-validation [zhang1993model]), for which the statistically optimal values may still depend on the intrinsic dimension of $\mathcal{D}$ or on the structure of $\bm{\theta}^{\star}$ [cawley2010over, raskutti2011minimax, negahban2011estimation].
Alternatively, a recent widely practiced yet theoretically underexplored approach is to over-parameterize the model, that is, to craft a sufficiently large region $\mathcal{D}^{\prime}$ that safely encompasses $\mathcal{D}$. This is typically achieved by overestimating the dimension $\dim(\mathcal{D})=k$ with $k^{\prime}\geq k$. However, directly solving $\min_{\bm{\theta}\in\mathcal{D}^{\prime}}L(\bm{\theta})$ often leads to overfitting, as the optimal solution might be in $\mathcal{D}^{\prime}\setminus\mathcal{D}$, resulting in poor statistical performance. To mitigate this issue, a common strategy is to introduce a reparameterization map $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{n}$ whose range covers $\mathcal{D}^{\prime}$, i.e., $\mathcal{D}^{\prime}\subseteq\{\varphi(\mathbf{x}):\mathbf{x}\in\mathbb{R}^{d}\}$. Therefore, the over-parameterization procedure transforms problem (1) into:
$\underbrace{\min_{\bm{\theta}\in\mathcal{D}}\ L(\bm{\theta})}_{\text{\bf Exactly-parameterized}}\stackrel{{\scriptstyle\text{\bf Over-parameterization}}}{{\implies}}\underbrace{\min_{\bm{\theta}\in\mathcal{D}^{\prime}}\ L(\bm{\theta})\quad\stackrel{{\scriptstyle\text{\bf Reparameterization}}}{{\implies}}\quad\min_{\mathbf{x}\in\mathbb{R}^{d}}\ f(\mathbf{x}):=L(\varphi(\mathbf{x}))}_{\text{\bf Over-parameterized}}.$ (2)
Over-parameterized models have become prominent in sparse recovery, where the reparameterization map is often defined as $\varphi:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ with $\varphi(\mathbf{u},\mathbf{v})=\mathbf{u}\odot\mathbf{v}$ [zhao2019implicit] or $\varphi(\mathbf{u},\mathbf{v})=\mathbf{u}\odot\mathbf{u}-\mathbf{v}\odot\mathbf{v}$ [vaskevicius2019implicit]. Here, the notation $\odot$ denotes the element-wise product. Additionally, over-parameterized models are widely used in low-rank matrix and tensor models when the true rank $r$ is unknown. In such scenarios, the true rank $r$ is often over-estimated with a search rank $r^{\prime}\geq r$, and the reparameterization map is defined as $\varphi:\mathbb{R}^{n\times r^{\prime}}\to\mathbb{R}^{n\times n}$ with $\varphi(\mathbf{X})=\mathbf{X}\mathbf{X}^{\top}$ for PSD matrices [chi2019nonconvex].
@@
Despite these challenges, recent research has shown that, in a wide range of over-parameterized learning problems with proper reparameterization maps, gradient-based algorithms exhibit an implicit regularization that steers them away from solutions outside $\mathcal{D}$, favoring those in the low-dimensional feasible region $\mathcal{D}$. For example, in sparse recovery, GD initialized at a small point and applied to the composite function $f(\mathbf{u},\mathbf{v})=L(\mathbf{u}\odot\mathbf{u}-\mathbf{v}\odot\mathbf{v})$ converges to the true $k$-sparse parameter $\bm{\theta}^{\star}$ [vaskevicius2019implicit, woodworth2020kernel], even though denser solutions may achieve comparable or even lower loss values. Remarkably, this guarantee holds even when the sparsity level $k$ is unknown. A related phenomenon arises under the reparameterization $\varphi(\mathbf{u},\mathbf{v})=\mathbf{u}\odot\mathbf{v}$ [zhao2019implicit, ma2022blessing]. Analogously, in low-rank matrix recovery, GD with a small initialization applied to $f(\mathbf{X})=L(\mathbf{X}\mathbf{X}^{\top})$ converges to the true rank-$r$ matrix $\mathbf{\Theta}^{\star}$, even when $r$ is unknown and severely overestimated (up to $r^{\prime}=n$[li2018algorithmic, stoger2021small, ma2023global]. Extensions of these results have also been established for low-rank tensor recovery [anandkumar2017analyzing], and various types of nonlinear neural networks [ma2023behind].
Despite these intriguing and promising findings, the theoretical foundations of the implicit regularization induced by GD in over-parameterized settings remain largely underexplored. In this paper, we aim to address this gap by presenting unified conditions that explain the implicit regularization in GD, potentially extending its applicability to broader classes of problems.
1.1 Our Goal
Our objective is to establish general conditions under which GD converges to a second-order stationary point (SOSP) of $f$ within the set $\varphi^{-1}(\mathcal{D})=\{\mathbf{x}\in\mathbb{R}^{d}:\varphi(\mathbf{x})\in\mathcal{D}\}$. For brevity, we refer to $\varphi^{-1}(\mathcal{D})$ as the implicit region and denote it by $\mathcal{M}:=\varphi^{-1}(\mathcal{D})$. Accordingly, an SOSP of $f$ within $\mathcal{M}$ will be called an $\mathcal{M}$-SOSP and is formally defined as:
$\text{$\mathbf{x}$ is $\mathcal{M}$-SOSP}\qquad\iff\qquad\mathbf{x}\in\mathcal{M},\ \ \nabla f(\mathbf{x})=0,\ \ \nabla^{2}f(\mathbf{x})\succeq 0.$ (3)
Our interest in an $\mathcal{M}$-SOSP stems from the key observation that, for generic functions, any $\mathcal{M}$-SOSP corresponds to a local minimum of $f$ within $\mathcal{M}$ [davis2019active, Theorem 2.9]. By construction of $\mathcal{M}$, any solution $\mathbf{x}\in\mathcal{M}$ maps to a parameter $\bm{\theta}=\varphi(\mathbf{x})\in\mathcal{D}$, which, by definition of $\mathcal{D}$, is low-dimensional and satisfies the feasibility constraints. In general, however, a local minimum $\mathbf{x}\in\mathcal{M}$ of $f$ need not map to a local minimum $\varphi(\mathbf{x})\in\mathcal{D}$ of $L$; indeed, it is possible for a local minimum of $f$ to map under $\varphi$ to a non-stationary point of $L$ [levin2025effect, Example 3.7]. Nevertheless, [levin2025effect, Theorem 2.3] shows that local minima of $f$ do map to local minima of $L$ under the mild assumption that $\varphi$ is an open map. For the canonical reparameterizations used in sparse recovery and low-rank matrix recovery, the openness of $\varphi$ can be verified directly [levin2025effect, Section 2]. Therefore, for these instances, any local minimum $\mathbf{x}\in\mathcal{M}$ of $f$ maps to a local minimum $\varphi(\mathbf{x})\in\mathcal{D}$ of $L$. Motivated by this connection, our goal is to address the following question:
Under what conditions on the implicit region $\mathcal{M}$, the objective function $f$, and the initialization $\mathbf{x}_{0}$ does GD converge to an $\mathcal{M}$-SOSP of $f$?
In this work, we identify three key conditions that enable the convergence of GD to an $\mathcal{M}$-SOSP: (i) initialization sufficiently close to the implicit region $\mathcal{M}$; (ii) efficient escape from strict saddle points (SSPs); and (iii) preservation of proximity to $\mathcal{M}$ throughout the iterations. Next, we explain these conditions and outline the main challenges associated with ensuring that they hold.
1.2 Major Challenges
Somewhat surprisingly, even if $\mathcal{D}$ (and consequently $\mathcal{M}$) are unknown a priori, it is still possible to select an initialization $\mathbf{x}_{0}$ within the proximity of $\mathcal{M}$ through an appropriate choice of $\varphi$. For instance, recall that in low-rank matrix recovery, the feasible region is the set of all PSD matrices with rank at most $r$, and the reparameterization is given by $\varphi(\mathbf{X})=\mathbf{X}\mathbf{X}^{\top}$. In this case, the zero matrix belongs to $\mathcal{M}=\varphi^{-1}(\mathcal{D})$ for every $r\leq n$. Similarly, in sparse recovery, the feasible region is the set of all vectors with sparsity level at most $k$, with reparameterizations $\varphi(\mathbf{u},\mathbf{v})=\mathbf{u}\odot\mathbf{u}-\mathbf{v}\odot\mathbf{v}$ or $\varphi(\mathbf{u},\mathbf{v})=\mathbf{u}\odot\mathbf{v}$. Here too, a pair of all-zero vectors lie in $\varphi^{-1}(\mathcal{D})$ for any $k\leq n$. As we will discuss later, analogous initializations can also be constructed for more general models with sparse basis function decompositions. Thus, initializing GD at these points satisfies the first condition.
@@ 1.2 Major Challenges
Now, we turn to the second condition, that is, the efficient escape from SSPs. It has been shown that various gradient-based algorithms, such as perturbed gradient descent (PGD) [ge2015escaping, jin2017escape] and stochastic gradient descent (SGD) [fang2019sharp], can escape SSPs and converge to an approximate SOSP as effectively as their second-order counterparts. To illustrate the underlying success mechanisms of these algorithms, let us consider PGD. Starting from an initial point $\mathbf{x}_{0}$, PGD iteratively updates the solution using the rule $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\nabla f(\mathbf{x}_{t})+\bm{\xi}_{t}$, where $\bm{\xi}_{t}$ represents an isotropic perturbation uniformly drawn from the Euclidean ball with radius $\gamma$ centered at zero when $\left\lVert\nabla f(\mathbf{x}_{t})\right\rVert$ is sufficiently small; otherwise, $\bm{\xi}_{t}$ is set to zero [jin2017escape, Algorithm 1]. The reasoning behind this episodic perturbation is natural: when $\left\lVert\nabla f(\mathbf{x}_{t})\right\rVert$ is small, the iterate $\mathbf{x}_{t}$ is either near an SSP or an SOSP. Unlike SOSPs, SSPs are unstable fixed points of GD. Consequently, an isotropic random perturbation of sufficient magnitude enables GD to escape from SSPs efficiently [jin2017escape, Theorem 3].
@@
Figure 1: PGD applied to matrix sensing converges under exact parameterization but fails to converge with over-parameterization. In this example, the true matrix $\mathbf{\Theta}^{\star}$ is a $20\times 20$ PSD matrix with rank $r=3$. The search rank is set to $r^{\prime}=3$ and $r^{\prime}=4$ for exact- and over-parameterized settings.
@@
Figure 2: When perturbation is added to the gradient, it contributes to two directions: deviation direction, which steers the iterations away from the implicit region $\mathcal{M}$, and the negative Hessian direction, which enables an escape from SSPs.
@@
Unfortunately, in over-parameterized settings, guaranteed escape from SSPs may come at the cost of significant deviation from the implicit region $\mathcal{M}$, thereby violating the third condition and preventing PGD from converging to an $\mathcal{M}$-SOSP. In fact, even mild over-parameterization can disrupt the convergence of PGD. Figure 2 illustrates this phenomenon on an instance of matrix sensing (introduced in Section 1). When the rank $r$ of the target matrix is known, PGD initialized at random and applied to $f(\mathbf{X})=L(\mathbf{X}\mathbf{X}^{\top})$ with $\mathbf{X}\in\mathbb{R}^{n\times r}$ converges to the ground-truth rank-$r$ matrix. By contrast, even a slight over-parameterization with a search rank of $r^{\prime}=r+1$ causes PGD to deviate significantly from $\mathcal{M}$ and converge to a solution of rank strictly larger than $r$, thereby incurring substantial error.
@@
The failure of PGD to converge to an $\mathcal{M}$-SOSP in the over-parameterized regime is precisely due to the impact of perturbations, as shown in Figure 2. When an isotropic random perturbation is applied, it inevitably affects two directions: a direction spanned by the eigenvectors corresponding to the negative eigenvalues of the Hessian, which pushes the iterate away from the SSP (shown as $(a)$), and another direction that causes the iterate to deviate away from the implicit region $\mathcal{M}$ (shown as $(b)$). When the perturbation is both isotropic and large, these directions become comparable, causing the PGD iterates to escape the saddle point but simultaneously drift away from $\mathcal{M}$, ultimately leading to the failure of the algorithm.
The central premise of this paper is that the second and third conditions can be met simultaneously: SSPs can be escaped efficiently via infinitesimal perturbations, while closeness to $\mathcal{M}$ is maintained by quantifying the deviation rate of $f$ from $\mathcal{M}$. We show that GD exhibits implicit regularization toward an $\mathcal{M}$-SOSP when initialized near $\mathcal{M}$ and equipped with both infinitesimal perturbations and a small deviation rate. Moreover, while the guaranteed escape from SSPs via infinitesimal perturbations is largely independent of $\mathcal{D}$ and the specific choice of $\varphi$, ensuring a small deviation rate depends crucially on $\mathcal{D}$ and the choice of $\varphi$. We establish that these conditions hold provably for the class of over-parameterized matrix sensing problems, and empirically for broader classes of problems.